### v1

# Chemical Language Model Linker: blending text and molecules with modular adapters

Abstract
The development of large language models and multi-modal models has enabled the appealing idea of generating novel molecules from text descriptions. Generative modeling would shift the paradigm from relying on large-scale chemical screening to find molecules with desired properties to directly generating those molecules. However, multi-modal models combining text and molecules are often trained from scratch, without leveraging existing high-quality pretrained models. That approach consumes more computational resources and prohibits model scaling. In contrast, we propose a lightweight adapter-based strategy named Chemical Language Model Linker (ChemLML). ChemLML blends the two single domain models and obtains conditional molecular generation from text descriptions while still operating in the specialized embedding spaces of the molecular domain. ChemLML can tailor diverse pretrained text models for molecule generation by training relatively few adapter parameters. We find that the choice of molecular representation used within ChemLML, SMILES versus SELFIES, has a strong influence on conditional molecular generation performance. SMILES is often preferable despite not guaranteeing valid molecules. We raise issues in using the large PubChem dataset of molecules and their associated descriptions for evaluating molecule generation and provide a filtered version of the dataset as a generation test set. To demonstrate how ChemLML could be used in practice, we generate candidate protein inhibitors and use docking to assess their quality.
1 Introduction
Machine learning methods have emerged as powerful tools in biology and chemistry (Ching et al., 2018). These methods have been applied extensively across a range of small molecule and protein problems, from molecular property prediction (Gilmer et al., 2017) and molecule interaction prediction (Deng et al., 2020) to protein structure prediction (Krishna et al., 2024) and inverse folding design (Zheng et al., 2023).
Generative machine learning models are a particularly attractive area of research as an alternative to traditional experimental and computational chemical screening. Directly generating molecules with desired properties can avoid performing high-throughput chemical screens and large-scale virtual screens on molecule databases (Alnammi et al., 2023). By learning the underlying chemical patterns and structures from vast datasets, generative models such as generative adversarial networks (Kadurin et al., 2017), variational autoencoders (Jin et al., 2018), and autoregressive models (Bagal et al., 2021) have enabled the automated design of molecules with desired properties. Recent efforts in generative chemical modeling include controllable cross-modality generation. These methods include contrastive learning, multi-task training, and finetuning of large language models (LLMs).
LLMs have dramatically reshaped the landscape of natural language processing (NLP), demonstrating remarkable capabilities in generating coherent and contextually appropriate text across diverse domains. Built on transformer architectures and trained on broad corpora, LLMs achieve a nuanced representation of language patterns, syntax, and semantics. This advanced representation enables them to perform a wide array of tasks, from simple text completion to complex question answering and summarizing. A critical aspect of their development has been the observed scaling laws (Kaplan et al., 2020), which suggest that the performance of these models improves predictably with an increase in model size and data volume. With LLMs’ achievements across different domains, researchers are increasingly exploring their potential to address challenges in the fields of biology and chemistry (Zhang et al., 2024).
Although general purpose LLMs show some capability to operate with molecules, it is limited (Christofidellis et al., 2023; Microsoft Research AI4Science and Microsoft Azure Quantum, 2023; Mirza et al., 2024). A natural question is how to construct a multi-modal model that operates on natural language and chemicals. Previous strategies include aligning the embedding spaces between text and molecules through contrastive learning, T5-based (Raffel et al., 2020) text-to-molecule translation, training a single model on multiple molecule-related tasks, and finetuning LLMs on molecule-specific corpora. However, all these methods have varying limitations. The T5 model requires additional pretraining and cannot apply previously pretrained models on related areas. Most notably, prior work typically combines a single natural language model with a single molecular model without exploring the interactions between models and molecular representations.
To address these issues, we introduce a flexible approach for merging arbitrary text-based language models with molecule generators named Chemical Language Model Linker (ChemLML). A chemical linker is a molecular structure that connects two or more molecules or functional groups via stable chemical bonds, widely used in drug development (for example, proteolysis targeting chimeras (Sakamoto et al., 2001) and fragment-based drug design (Grenier et al., 2023)), biomolecular research, and material science. We borrow the idea of “linker”, and our “linker” metaphorically connects natural and molecular languages through adapters instead of linking different molecules or functional groups. Despite the name, ChemLML is not specifically designed to generate linker molecules (Imrie et al., 2020; Guan et al., 2023; Igashov et al., 2024). ChemLML makes full use of pretrained models for text and molecules, which can greatly reduce the training time. The main contributions are:
  1. 1.
    We present ChemLML, which requires significantly fewer trainable parameters compared to models of the same scale for text-guided molecule design tasks, achieving strong performance by training adapters only.
  2. 2.
    ChemLML can flexibly combine multiple types of pretrained text and molecule architectures.
  3. 3.
    Our case study suggests that ChemLML may be able to generate candidate inhibitors of protein targets from text descriptions alone based on its docking performance.
2 Related Work
Contrastive learning Contrastive learning has emerged as a powerful technique for learning rich, meaningful representations in multiple domains. SimCLR (Chen et al., 2020) performs contrastive learning between different augmented views of the same image to learn visual representations. MolCLR (Wang et al., 2022) extends a similar method to graph neural networks on molecule graphs. CLIP (Radford et al., 2021) introduces a novel approach to contrastive learning that unifies language and vision modalities. The model is trained to predict which images are paired with which texts. MoleculeSTM (Liu et al., 2023a) trains a multi-modal molecule-text model by learning a joint representation of molecules’ chemical structures and textual descriptions via contrastive learning. MoleculeSTM is able to perform structure-text retrieval and molecule editing tasks.
Multi-task learning Multi-task learning is applied to molecular tasks because it enables models to leverage shared information across tasks, often improving generalization and performance by learning common patterns. The origins of some multi-modal text-and-chemical models came from NLP. For instance, MT-DNN (Liu et al., 2019) leverages the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) architecture as a shared encoder and simultaneously learns across multiple NLP tasks. By adding task-specific layers on top of BERT, MT-DNN integrates the advantages of shared knowledge to improve generalization and performance on those tasks. The T5 model (Raffel et al., 2020) treats all NLP tasks as text-to-text problems, enabling unified, scalable multi-task learning with a single model. The model is pretrained and finetuned on diverse tasks for enhanced performance and efficiency. MolT5 (Edwards et al., 2022) uses a T5 model to perform molecule-to-text and text-to-molecule tasks at the same time. Text+ChemT5 (Christofidellis et al., 2023) further extends the number of molecule-related tasks.
Chemical language models Chemical language models are models that are trained on string representations of molecules (Grisoni, 2023). They primarily fall into two categories of architectures: BERT and generative pretrained transformer (GPT). ChemBERTa (Chithrananda et al., 2020) is a specialized adaptation of the BERT model, tailored for chemistry. It is pretrained with masked language modeling and finetuned for molecule property prediction. MolGPT is a GPT model tailored for molecular generation tasks that is trained by predicting the next token. For a more comprehensive understanding of specific models, see recent reviews (Zhang et al., 2024; Liao et al., 2024; Ramos et al., 2024).
Text-guided molecule generation Text2Mol (Edwards et al., 2021) proposes the problem of generating molecules from their text descriptions, which is framed as a molecule retrieval task. Instead of simply retrieving the most similar molecules or text, MolT5 (Edwards et al., 2022) trains a T5 model that can perform text-guided molecule generation and molecule captioning tasks at the same time. Text+Chem T5 (Christofidellis et al., 2023) further trains a T5 model on multiple tasks related to molecules and improves the model’s capability on text-guided molecule generation. nach0 (Livne et al., 2024) is a multi-modal model also based on the T5 architecture that is pretrained on PubMed abstracts, patent descriptions, and ZINC chemicals. MoMu (Su et al., 2022) uses contrastive learning to bridge a graph-based molecule encoder and a natural language encoder to perform multiple molecular tasks. TGM-DLM (Gong et al., 2024) leverages diffusion models to address the limitations of autoregressive methods in molecule generation. MolXPT (Liu et al., 2023c) includes wrapped sentences in its pretraining dataset, which are sentences from scientific text that contain molecules that are detected and replaced with the matching SMILES string. MolFM (Luo et al., 2023) includes a knowledge graph over drugs, proteins, and diseases in its multi-modal model. AMAN (Zhao et al., 2024a) introduces a graph transformer for the molecule representation and trains using a combination of adversarial loss and a triplet loss that includes negative samples. ChemDFM (Zhao et al., 2024b) performs chemistry domain pretraining and instruction tuning on multiple molecule-related tasks with Llama2, developing an dialogue-based chemical LLM. Mol-Instruction (Fang et al., 2023a) prepares datasets for instruction tuning LLMs, which includes text-guided molecule generation instructions. MolReGPT (Li et al., 2024) uses in-context learning with ChatGPT and Llama 2. In-context learning with ChatGPT was also used to generate synthetic training data for molecule generation (Chen et al., 2024). MoleculeSTM (Liu et al., 2023a) can perform molecule captioning, text-based molecule generation, and text-based molecule editing. A comprehensive review presents additional approaches Pei et al. (2024). The recent Language + Molecules Workshop (Edwards et al., 2024), whose dataset includes biomedical and other chemical properties, also inspired new methods.
3 Methods
3.1 Task definition
We define the task as generating molecules based on a text description such that the molecules match the description. The text may contain information about the desired physical, functional, or chemical properties of the molecule.
3.2 Preliminaries
Simplified molecular input line-entry system (SMILES) SMILES is a notation used to represent molecules as strings (Krenn et al., 2020). SMILES strings describe atomic composition, connectivity, and stereochemistry using simple rules. For example, the SMILES of 4-Methylphenol is Cc1ccc(O)cc1.
Self-referencing embedded strings (SELFIES) SELFIES is an alternative representation for molecules designed to overcome some limitations of SMILES (Krenn et al., 2020). It ensures the generation of syntactically and semantically valid molecular structures. This is achieved through a self-referencing system that encodes rules and constraints within the string itself, making SELFIES particularly useful for applications in generative models. The SELFIES of 4-Methylphenol is [C][C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][O]. It is much longer than the corresponding SMILES format.
3.3 ChemLML model architecture
Text pretrained models We used three pretrained text LLMs that are specialized for scientific text. SciBERT (Beltagy et al., 2019) is a lightweight model with 110M parameters. Galactica (Taylor et al., 2022) was reported to have the best performance in a chemistry-related evaluation (Mirza et al., 2024). Due to GPU memory constraints, we only used Galactica-125M, 1.3B, and 6.7B. We also used the T5 encoder from Text+ChemT5 (Christofidellis et al., 2023).
Molecule pretrained models We focused on pretrained models that use SMILES or SELFIES string representations. MolGPT (Bagal et al., 2021) is a GPT model. MolGen (Fang et al., 2023b) is a BART-style model that has two stages of pretraining. During the first stage, it randomly masks some of the SELFIES tokens, encodes the corrupted SELFIES using a bidirectional model, calculates the likelihood of SELFIES string with a left-to-right autoregressive decoder, and calculates the reconstruction loss. In the second stage, it introduces the domain-agnostic molecular prefix as a domain instructor to facilitate the transfer of knowledge across diverse domains. We only use the MolGen autoregressive decoder. MolXPT (Liu et al., 2023c) is another GPT-style model, which unifies text and molecules during pretraining. It replaces the molecule name with a SMILES string. We only use the molecule decoding capability of MolXPT by feeding the start-of-molecule token at the beginning during decoding.
Complete architecture We use the text pretrained model as a text encoder and the molecule pretrained model as a molecule decoder. We add a chemical adapter to the last layer of the molecule decoder that takes both text and molecule embeddings as input and outputs a molecule embedding. A similar architecture was applied in the protein domain with LM-Design (Zheng et al., 2023) and ProtT3 (Liu et al., 2024). The model architecture is shown in Fig. 1.
Figure 1: The ChemLML model architecture. We use a pretrained text model as the text encoder and a pretrained molecule model as the molecule decoder. A single modular chemical adapter is added to the last layer of the molecule decoder. It takes the text and molecule embeddings as input and produces a refined molecule embedding under the guidance of the text embedding. The fire and snowflake icons indicate the parameters are trainable or frozen. FFN: feed-forward network, LM: language modeling, BOS: beginning of sequence, EOS: end of sequence.
3.4 ChemLML training and inference
Adapter module We use a cross-attention mechanism to construct the adapter between the text embedding and molecule embedding. Let $T=\{t_{1},t_{2},...,t_{m}\}$ be the set of text embeddings and $S=\{s_{1},s_{2},...,s_{n}\}$ be the set of molecule token embeddings. In cross-attention, the text embedding will be projected to the same dimension as the molecule embedding, denoted as $T^{\prime}=\{t^{\prime}_{1},t^{\prime}_{2},...,t^{\prime}_{m}\}$. We can obtain the cross-attention matrix $a$ with size $m\times n$. For $i=1,2,...,m$ and $j=1,2,...,n$, the attention weights can be calculated as $a_{i,j}=\frac{\exp(t_{i}^{\top}s_{j}^{\prime})}{\sum_{k=1}^{m}\exp(t_{i}^{\top}s_{k}^{\prime})}$ and the molecule embedding will be updated as $s_{i}^{\prime}=\sum_{j=1}^{m}a_{i,j}s_{j}$. We input the new molecule embedding into the language model head of the molecule pretrained model. The updated embeddings still reside within the embedding space of the original pretrained molecule model. The difference is that these molecules are now aligned with the text embeddings, thereby inheriting the properties described by the text.
During the training process, we use textual descriptions as input, which are first fed into the text LLM model to obtain their embeddings. Subsequently, we use the adapter to perform the cross-attention operation between the text embeddings and molecule embeddings from the last layer of the molecule pretrained model to capture the relationship between the textual and molecular information. We adopt the teacher-forcing strategy during training, using the ground truth token as the input for next step rather than using the model’s own output from the last step.
During the inference stage, we employ an autoregressive approach to generate molecules. Specifically, the model generates the molecule sequence in a step-by-step manner, where each step predicts the next sequence token based on the previously generated sequence and the text embedding. This process can be expressed as:
$y_{t}\sim P(y|y_{<t},T;\theta)$ (1)
where $y_{t}$ is the token generated at time step $t$, $y_{<t}$ denotes the sequence generated prior to time step $t$, $T$ is the given text embedding, and $\theta$ represents the model parameters.
3.5 Metrics
We intend generated molecules to exhibit specific properties. Following the molecular similarity principle (Duran-Frigola et al., 2020), this would result in the generated molecules having structures similar to the ground truth molecules, thereby achieving a high degree of similarity and low diversity during evaluations. Therefore, we primarily focus on various similarity metrics and calculate the similarity between generated molecules and ground truth molecules. We use MACCS fingerprints, RDKit fingerprints (Landrum, 2018), and Morgan fingerprints as different molecule representations that produce different similarity scores. The similarities between these fingerprints are calculated with Tanimoto similarity. We discuss other metrics not used here in the Supplementary Methods.
SMILES-based generation will generate syntactically invalid molecules sometimes, so we also include validity as a metric. Even though SELFIES is designed to include 100% validity, the SELFIES-based model might fail to generate a valid molecule, such as samples with an [EOS] token at the beginning. The validity metric only evaluates syntactic correctness without accounting for the molecule’s 3D structure.
3.6 Datasets
Pretraining datasets The pretrained molecule generation models were trained on the following datasets:
  • Selected ZINC-15: Chemformer (Irwin et al., 2022) selected approximately 100M molecules from 1.5B available molecules from ZINC-15 (Sterling and Irwin, 2015) with the following constraints: reactivity set to reactive, purchasability set to annotated, molecular weight $\leq 500$ Daltons, and LogP $\leq 5$. MolGen uses this dataset during the first stage of pretraining.
  • NPASS: Natural Product Activity & Species Source Database (NPASS) (Zeng et al., 2018) is a natural product domains database, which is used in the second pretraining stage of MolGen.
  • MOSES: MOSES (Polykovskiy et al., 2020) is a cleaned version of ZINC, containing 1.9M molecules. It removes the atoms besides C, N, S, O, F, Cl, Br, H and cycles longer than 8 atoms. It does not contain chirality information. Models pretrained on the MOSES dataset, such as MolGPT, are unable to process SMILES strings that include atoms beyond this dataset’s scope or SMILES strings that contain chirality information.
  • PubChem and PubMed: MolXPT trains on the titles and abstracts from 30M PubMed entries, SMILES strings from 30M chemicals from PubChem (Kim et al., 2023), and 8M sequences in which chemical names are replaces with SMILES.
We do not train directly on these datasets. However, information about the training data of the molecule generation models that ChemLML builds upon is helpful for understanding its molecule generation performance.
Molecule description datasets We use two different molecule description datasets. We train and test all the models on ChEBI-20 and use PubChem as an additional test set. We use the training set of ChEBI-20 for training. Then, we tune the hyperparameters based on the convergence status on the ChEBI-20 training set and the performance on the ChEBI-20 validation set. Finally, we test on the ChEBI-20 and PubChem test sets.
  • ChEBI-20: The Chemical Entities of Biological Interest (ChEBI) database (Hastings et al., 2016) is a dictionary of small molecules that includes text descriptions. Edwards et al. (2021) created the ChEBI-20 dataset by combining compounds from PubChem with ChEBI annotations and retaining the 33,010 compound-description pairs for which the description had more than 20 words. The ChEBI-20 training, validation, and test set splits of 80%, 10%, and 10% are also from Edwards et al. (2021).
  • PubChem: In our evaluations, we assume that each text description is informative and associated with one unique molecule. Many examples in the complete PubChem dataset (Kim et al., 2023) do not meet that assumption, so we apply several types of filters on the molecules and text descriptions. In short, we select descriptions greater than 30 words and remove one-to-many description-molecule pairs. In this way, we remove ambiguous instances and improve the quality of our evaluations. After filtering, we remove the intersection of molecules between PubChem and ChEBI-20 dataset and obtain a dataset that has 11,563 examples. We randomly sample 3,000 examples from this dataset and use it as an additional test set for models trained on ChEBI-20. To demonstrate the impact of this filtering, we also report results on the unfiltered PubChem dataset with 2,576 molecule descriptions, which we refer to as PubChem-unfiltered. More details of PubChem dataset filtering and processing are discussed in Supplementary Methods.
3.7 Baseline methods
T5 The T5 model comes from the MolT5 paper (Edwards et al., 2022). It is trained on the ChEBI-20 dataset.
MolT5 MolT5 loads the T5.1.1 checkpoint and is pretrained using the replace corrupted spans objective. During each pretraining step, a minibatch comprising both natural language sequences and SMILES sequences is sampled and certain words are randomly selected for corruption. The task is to predict the spans of tokens that were corrupted. In the finetuning stage, the model is trained on ChEBI-20.
MolXPT MolXPT is pretrained on PubMed and PubChem data. Then, it is finetuned on ChEBI-20 dataset.
TGM-DLM TGM-DLM (Gong et al., 2024) uses the text embedding to guide a diffusion language model to generate molecules. It employs a two-phase diffusion process, first generating a molecule based on the text description followed by correcting invalid SMILES strings.
Text+ChemT5 Text+ChemT5 is a multi-task, multi-domain model for natural and chemical language. It is trained on the four tasks listed in the Supplementary Methods.
T5 encoder+MolXPT MLP This model concatenates the T5 text and MolXPT molecule embeddings and passes them through a two-layer multilayer perceptron (MLP). It serves as an ablation for ChemLML to demonstrate the value of the full ChemLML model over its individual components.
3.8 Training and evaluation
We evaluate the model’s performance on text-based molecule generation. For MolT5, Text+ChemT5, MolXPT, SciBERT, and different scales of MolGen, we use their checkpoints on HuggingFace. For the Text+ChemT5 encoder, we add the prompt "Write in SMILES the described molecule:", which is used in the corresponding task in Text+ChemT5.
For ChemLML, we use SciBERT, the encoder from Text+ChemT5 and different scales of Galactica as the text encoders and combine them with MolGPT, MolXPT, MolGen, and MolGen-7B as the molecule decoders. The weights of molecule decoder are always frozen. The weights of the adapter are trainable. The weights of the text encoder can be either trainable or frozen. We run experiments on two datasets, ChEBI-20 and PubChem.
For ChEBI-20, we report the results from the MolT5, Text+ChemT5, MolXPT, and TGM-DLM papers. T5 is included in the MolT5’s paper.
MolGPT can not be finetuned in ChEBI-20 because it is pretrained on MOSES dataset so its tokenizer can not deal with chirality information. We use MolGPT to compare SMILES- and SELFIES-based models. First, we use RDKit to remove stereochemistry information from all the molecules. Then, we remove all the molecules that fail to be tokenized by MolGPT’s tokenizer. Lastly, we transform all the SMILES to SELFIES. This reduces the ChEBI-20 training set from 26,407 to 15,899 instances, validation set from 3,301 to 1,961 instances, and test set from 3,300 to 2,032 instances. We build the MolGPT model with 12 layers, 8 attention heads, and 256 as the embedding dimension. The only differences between MolGPT SMILES and SELFIES models are the output dimensions, depending on the vocabulary lists of each method.
We use the Noam Optimizer with 4,000 warm up steps for training. For the sampling in the test set, we use multinomial sampling with random seed 42 for all methods.
3.9 Docking case study setting
We filtered the PubChem dataset to identify text descriptions that pertain to molecules that inhibit specific protein targets by initially retaining only descriptions with the string "inhibit". Then, we split all remaining instances into low (0.15 to 0.3), medium (0.3 to 0.5), and high (0.5 to 0.9) similarity based on the Morgan fingerprint similarity between the ground truth and generated molecule from the ChemLML (T5 encoder+MolGen) model. We selected one example per similarity bin (Table 7.4) as described in the Supplementary Methods.
After we generated the first molecule with the temperature set to 1 and random seed 42, we set the temperature to 1.5 and changed the random seed in order to sample different molecules. We canonicalized the generated SMILES with RDKit and removed duplicates. We repeated this process until we obtained 99 additional unique and valid SMILES, resulting in a total of 100 ChemLML-generated molecules per target. The SMILES were then used to build 3D conformers with OpenEye’s Omega2, and only a subset produced valid 3D structures for docking (Supplementary Methods). For ChemLML, we used the T5 encoder+MolXPT and T5 encoder+MolGen settings. We omit "encoder" for brevity in the docking experiments. We performed the same molecule generation process with the MolT5 and Text+ChemT5 baselines.
For each target protein, we docked the ground truth molecule, generated molecules, and control molecules from two types of background distributions: FDA-approved compounds and ChemLML T5+MolGen-generated molecules from randomly-sampled descriptions. A consensus docking score was computed for each molecule that combines results from four docking programs (Supplementary Methods). We hypothesized that the ground truth molecule and generated molecules resembling this ground truth molecule would score more favorably on their respective protein targets than background molecules.
3.10 Code and data availability
The ChemLML code and PubChem-filtered dataset are available from https://github.com/gitter-lab/ChemLML and archived at https://doi.org/10.5281/zenodo.13925649. The pretrained ChemLML models and PubChem-unfiltered dataset are available from https://zenodo.org/doi/10.5281/zenodo.11661517. The ChEBI-20 dataset is available from https://github.com/cnedwards/text2mol, and a copy is in the ChemLML GitHub repository.
4 Results
We first compare baseline and ChemLML models trained on ChEBI-20 text descriptions on the ChEBI-20 test set (Table 1). The best performing ChemLML model is the T5 encoder finetune+MolXPT version. With 114M trainable parameters, this model achieves 0.727 in Morgan FTS. It is not surprising to find that models that include the T5 encoder work the best because the T5 encoder has been trained on multiple tasks. The ChemLML combination T5 encoder+MolXPT performs better than the baseline MolT5 even though MolT5 has 52 times more trainable parameters.
Format Models Trainable/Total Params. Exact $\uparrow$ MACCS FTS $\uparrow$ RDK FTS $\uparrow$ Morgan FTS $\uparrow$ Validity $\uparrow$
Baseline methods SMILES T5 247M/247M 0.069 0.731 0.605 0.545 0.660
SMILES MolT5 248M/248M 0.081 0.721 0.588 0.529 0.772
SMILES MolXPT 350M/350M 0.215 0.859 0.757 0.667 0.983
SMILES TGM-DLM 180M/180M 0.242 0.854 0.739 0.688 0.871
SMILES Text+ChemT5 223M/223M 0.212 0.874 0.767 0.697 0.792
SMILES T5 encoder+MolXPT MLP  [^1] M/461M 0.227 0.820 0.713 0.636 0.980
ChemLML SELFIES SciBERT+MolGen 4.7M/317M 0.083 0.680 0.526 0.422 0.995
SELFIES Galactica 125M+MolGen 4.7M/ [^3] M 0.072 0.687 0.533 0.433 0.994
SELFIES T5 encoder+MolGen 4.7M/317M 0.165 0.758 0.616 0.527 0.995
SMILES T5 encoder+MolXPT 4.7M/464M 0.305 0.837 0.745 0.674 0.986
SELFIES Galactica 1.3B+MolGen 7.4M/1.53B 0.091 0.706 0.560 0.454 0.995
SELFIES Galactica 6.7B+MolGen 11.5M/6.87B 0.103 0.732 0.588 0.480 0.997
SELFIES Galactica 125M+MolGen 7B 56.7M/6.66B 0.134 0.732 0.608 0.522 0.995
SELFIES T5 encoder+MolGen 7B 56.7M/6.64B 0.209 0.777 0.666 0.581 0.995
SELFIES SciBERT finetune+MolGen 115M/317M 0.176 0.757 0.633 0.547 0.996
SELFIES T5 encoder finetune+MolGen 114M/317M 0.250 0.790 0.667 0.587 0.994
SMILES T5 encoder finetune+MolXPT 114M/464M 0.389 0.868 0.790 0.727 0.987
SELFIES Galactica 125M finetune+MolGen 130M/ [^3] M 0.143 0.743 0.606 0.510 0.995
SELFIES Galactica 125M finetue+MolGen 7B 182M/6.66B 0.170 0.759 0.640 0.554 0.995
SELFIES T5 encoder finetune+MolGen 7B 166M/6.64B 0.234 0.793 0.678 0.596 0.995
Table 1: Results of molecule generation on the ChEBI-20 test set. The T5, MolT5, MolXPT, TGM-DLM, and Text+ChemT5 results are copied from their respective papers. The ChemLML models are grouped into those that finetune the text model and those that do not. We bold the best model per metric within each model category: baseline, ChemLML without finetuning, and ChemLML with finetuning. FTS: Fingerprint Tanimoto similarity
4.1 Comparison among pretrained models used with ChemLML
When comparing the ChemLML models that use the MolGen model and do not finetune the text model, it is surprising that the cross-modal representation capability of similarly sized text models does not grow over time. The ChemLML variants SciBERT+MolGen and Galactica 125M+MolGen performed similarly. SciBERT was published in 2018 with 110M parameters and is trained on 3.3B tokens. Galactica 125M is trained on 106B tokens, around 32 times more than SciBERT. A possible explanation is that Galactica 125M is too small for the large training set. However, Galactica 1.3B also outperforms SciBERT by only a small margin. Even Galactica 6.7B, which is far larger than Galactica 125M, only yields a slight improvement. LLMs perform well when combined with MolGen 7B. ChemLML Galactica 125M+MolGen 7B achieves better results than MolT5.
We also consider the results from finetuning the pretrained language models along with the adapter. SciBERT outperforms Galactica 125M in this setting. Also, ChemLML SciBERT finetune+MolGen outperforms MolT5 with half the trainable parameters. The combination of the T5 encoder and MolGen is further strengthened by finetuning the T5 encoder. However, for the ChemLML T5 encoder models that use MolGen 7B, the improvement from finetuning the T5 encoder is marginal.
Based on the values in Table 1, it seems that ChemLML T5 encoder+MolGen does not outperform the baseline Text+ChemT5 on all similarity metrics at first glance. However, the results are influenced by the trade-off between similarity and validity and the selection of SMILES and SELFIES, which we expand upon below.
4.2 Molecule similarity and validity trade-off
There is a trade-off between generated and ground truth molecule similarity and the generated molecule validity (Table 1). T5 and MolT5 have the same model architecture. MolT5 has been pretrained on a large amount of data with both SMILES and natural language while T5 has not. At first glance, it is therefore surprising that T5 outperforms MolT5 in the similarity metrics. However, MolT5 generates 11 percentage points more valid molecules than T5, which means more molecules are counted in the similarity calculation. These molecules that are more difficult to generate may have lowered the overall similarity.
4.3 SMILES versus SELFIES representations
The choice of molecule representation, SMILES or SELFIES, is another factor that influences the similarity calculations. Skinnider (2024) recently showed that a molecule language model trained on SMILES had Fréchet ChemNet Distance (FCD) lower than 0.5, while one trained on SELFIES had FCD greater than 2 because invalid SMILES are low-likelihood samples and they improve performance on distribution-learning metrics like FCD. However, MolGen is trained on SELFIES, and there is not yet an equivalent model pretrained on SMILES that goes through multiple stages of pretraining. Therefore, we use MolGPT to compare SMILES and SELFIES directly.
Representation Trainable/Total Params Exact $\uparrow$ MACCS FTS $\uparrow$ RDK FTS $\uparrow$ Morgan FTS $\uparrow$ Validity $\uparrow$
SMILES 0.59M/120M 0.035 0.704 0.508 0.454 0.496
SELFIES 0.59M/120M 0.020 0.620 0.398 0.308 1.000
SELFIES (on valid SMILES) 0.59M/120M 0.021 0.616 0.400 0.303 1.000
Table 2: Comparison of MolGPT models trained on SMILES and SELFIES representations.
After MolGPT models are pretrained on the MOSES dataset, we train ChemLML T5 encoder+MolGPT models on the ChEBI-20 training dataset and evaluate the performance on the test set (Table 2). There is a large difference between the performance of SMILES and SELFIES versions of the same model. When we limit the molecules to the intersection of valid molecules between both methods to ensure the results are unaffected by the similarity and validity tradeoff we mention above, the SELFIES method still performs poorly. The SMILES method outperforms the SELFIES method by about 50% in the Morgan fingerprint similarity metric. Therefore the prior study’s conclusions about unconditional molecule generation (Skinnider, 2024) also hold for conditional molecule generation.
Based on SMILES substantially outperforming SELFIES in this analysis, we focus on comparing different pretrained models instead of comparing between SMILES- and SELFIES-based methods.
4.4 PubChem evaluation reveals pros and cons of LLMs
Because many of the PubChem molecule descriptions are overly ambiguous and generic, we prioritize the PubChem-filtered version of the dataset but also evaluate the PubChem-unfiltered dataset to assess how the filtering affects the results. On PubChem-filtered, LLMs without finetuning perform poorly, while the finetuned LLMs encoders show better results (Table 3). PubChem-filtered has a similar distribution as the ChEBI-20 dataset, so the ChemLML T5 encoder+MolXPT model still outperforms other methods that have a frozen text encoder. The ChemLML combination T5 encoder finetune+MolGen 7B does not perform well, potentially due to overfitting on the ChEBI-20 dataset. Another notable finding is that ChemLML generates a higher proportion of exactly matched molecules compared to the MolT5 and Text+ChemT5 baselines. We suggest that it is because MolGen and MolXPT have been pretrained on a large number of molecules, then ChemLML is able to retrieve those molecules accurately based on the text embedding.
Models Trainable/Total Params Exact $\uparrow$ MACCS FTS $\uparrow$ RDK FTS $\uparrow$ Morgan FTS $\uparrow$ Validity $\uparrow$
Baseline methods Text+ChemT5 223M/223M 0.028 0.682 0.584 0.495 0.696
MolT5 248M/248M 0.052 0.672 0.600 0.594 0.616
ChemLML T5 encoder+MolGen 4.7M/317M 0.165 0.588 0.477 0.376 0.967
T5 encoder+MolXPT 4.7M/464M 0.247 0.610 0.526 0.428 0.978
Galactica 125M+MolGen 4.7/ [^3] M 0.002 0.458 0.311 0.204 0.968
Galactica 1.3B+MolGen 7.4M/1.53B 0.004 0.489 0.332 0.221 0.958
T5 encoder+MolGen 7B 56.7M/6.64B 0.004 0.483 0.341 0.225 0.988
T5 encoder finetune+MolGen 114M/317M 0.215 0.598 0.505 0.404 0.973
T5 encoder finetune+MolXPT 114M/464M 0.286 0.624 0.543 0.453 0.972
Galactica 125M finetune+MolGen 130M/ [^3] M 0.191 0.566 0.469 0.372 0.978
T5 encoder finetune+MolGen 7B 166M/6.64B 0.008 0.532 0.388 0.265 0.966
Table 3: Results of molecule generation on the PubChem-filtered test set. SciBERT is excluded because its encoder cannot handle inputs exceeding 512 tokens. The ChemLML models are grouped into those that finetune the text model and those that do not, and we bold the best model in each of those categories.
We are cautious interpreting results from the PubChem-unfiltered dataset (Table S1) due to the abundance of generic descriptions (Figure S1), but they do illuminate difference among the ChemLML models. We observe that the performance of Galactica 125M surpasses both SciBERT and the T5 encoder when they are all used with frozen text encoders and the MolGen model, contrary to what was observed with the ChEBI-20 dataset. The large pretraining corpus on scientific texts might account for this. It could enable the models to perform better on unseen descriptions.
4.5 Docking case study
To demonstrate how ChemLML could be used in practice, we applied it to generate chemical inhibitors of protein drug targets acetylcholinesterase, inosine-5’-monophosphate dehydrogenase, and heat shock protein HSP 90-alpha (Table 7.4). We obtained the experimental structures of these proteins and used FRED (McGann, 2011), Gnina (McNutt et al., 2021), PLANTS (Korb et al., 2009), and rDock (Ruiz-Carmona et al., 2014) to dock the ground truth inhibitor (co-crystallized ligand) from PubChem, the ChemLML-generated candidate inhibitors, and candidate inhibitors generated by two baseline methods. In addition, we docked control molecules generated by ChemLML and an FDA-approved screening library that are not expected to bind the targets. We generated a single consensus docking score per compound (Figure 2).
Figure 2: Docking score distributions for the three protein targets in Table 7.4. Higher scores on the y-axis are better. Sample sizes for the docked compound sets are: FDA 2,956 for all targets, ChemLML_background (bg) 657 for all targets, ChemLML T5+MolGen 16/42/52, ChemLML T5+MolXPT 99/99/92, MolT5 97/98/96, and Text+ChemT5 83/96/86.
Docked ground truth and most generated molecules score poorly for the heat shock protein HSP 90-alpha, likely due to poor handling of macrocycles by docking programs (Table 7.4). Here, control compounds from the FDA and ChemLML background sets have median scores (0.00009 and 0.00043, respectively) similar to the ground truth molecule (0.00006) and the ChemLML T5 + MolGen generated ensemble median (0.00004). The ChemLML T5+MolXPT generated ensemble scores slightly higher with a median of 0.00115 but not substantially better than the random ChemLML background. Considering the poor docking score of the ground truth molecule, we consider the HSP 90-alpha target a poor system for comparative performance evaluation of the molecule generators. Also, we cannot draw conclusions about the quality of the 3D conformers built from the SMILES of the generated molecules.
For inosine-5’-monophosphate dehydrogenase, the ground truth molecule and ChemLML T5+MolGen ensemble score markedly better (0.00488 and median of 0.00594) than the control FDA and ChemLML background sets (medians of 0.00006 and 0.00049). The ChemLML T5+MolXPT ensemble, however, does not show more favorable scores with a median score near the backgrounds (0.00040). The difference in median scores could be related to the difference in valid 3D conformations produced by each of these ChemLML variants.
For acetylcholinesterase, the ChemLML T5+MolGen generated molecules have docking scores (median 0.00237) well above the background distributions for FDA and generated sets (medians of 0.00013 and 0.00064) but below the ground truth molecule (0.00873). The other generated ensembles—ChemLML T5+MolXPT, MolT5, and Text+ChemT5—do not perform better than the background sets. The first generated molecule from ChemLML T5+MolGen has a docking score (0.00482) above the background distributions, yet departs structurally from the ground truth molecule (Table 7.4), which is valuable in drug discovery. We also inspected whether the generated molecules satisfied the chemical structural properties in the text descriptions and their chemical structural diversity (Supplementary Results).
5 Discussion
This study proposes the ChemLML framework for using pretrained language and molecular models in text-guided molecule design. By reusing large-scale pretrained models, we enhance the flexibility of text-based molecule generation and reduce the training effort by supporting multiple types of frozen text encoders. In our case study generating candidate inhibitors for three protein targets from text, we find that for the two targets where docking performs well, ChemLML T5 encoder+MolGen generates molecules with docking scores that are better than ChemLML T5 encoder+MolXPT, two existing molecule generation algorithms, and two distributions of control molecules. For the inosine-5’-monophosphate dehydrogenase target, the median ChemLML docking score is even better than the ground truth inhibitor from PubChem.
When finetuning LLMs, we find it is easy to finetune Galactica 125M. However, it becomes harder to finetune the Galactica 1.3B model. Other finetuning methods such as Low-Rank Adaptation (Hu et al., 2022) may solve this problem. Also, we do not carry out experiments on Galactica 30B and 120B due to hardware and training technique limitations. Improving the LLM finetuning and finetuning larger LLMs are directions for future work.
A limitation of ChemLML is that it only focuses on molecule generation whereas related methods are multi-task. For instance, MolT5 can perform molecule captioning, and Text+ChemT5 can also conduct forward reaction and retrosynthesis prediction. Conceptually, the ChemLML framework is compatible with a multi-task setting and even more modalities including proteins (Liu et al., 2023b, 2024). This presents another avenue for future work.
Even though ChemLML performs well on the CheBI-20 and PubChem datasets, there is still a gap between these experiments and real world applications. The molecular descriptions in these datasets follow a relatively fixed format. We have not conducted user studies to assess how chemists’ general expectations for text-based molecule generation deviate from the types of descriptions in CheBI-20 and PubChem. In addition, a limitation of text-based molecule generators like ChemLML is that the text encoder can be sensitive to the prompt. The necessity to conduct evaluations on datasets like CheBI-20 and PubChem, where molecules have existing text descriptions, makes it challenging to evaluate these models’ robustness and generalizability to out-of-distribution data, both new text prompts and new molecules.
Even after our PubChem filtering, data quality issues in this dataset likely remain (Supplementary Methods). PubChem-unfiltered offers a large potential training dataset, and similar datasets have been used in prior work. We intentionally decided to not train models on the generic text descriptions in PubChem and caution others to think carefully about whether PubChem-unfiltered or PubChem-filtered is more appropriate for their training and evaluation goals.
Many molecule generation models, including ChemLML, still face limitations in generating fully plausible, semantically-valid molecules even when then are syntactically-valid. We observed that ChemLML-generated molecules in the SMILES or SELFIES format may fail to be transformed to 3D conformers for docking. In general, inspection of other molecule generation methods has uncovered unstable ring systems, reactive or toxic moieties, or geometric errors (Walters, 2024). ChemLML may be susceptible to these issues as well.
Finally, generating structurally similar molecules does not necessarily ensure that they will share the same properties as the target molecule. Many molecule descriptions pertain to the chemical structure of the molecules. In these cases, the assumption made by the chemical structure-based evaluations is most appropriate. For other physicochemical properties, different chemical structures could yield similar properties. In these cases, the selection of improved evaluated metrics remains another area for future exploration.
6 Conclusion
ChemLML introduces the strategy of using an adapter between text LLMs and molecule decoders for text-guided molecule design. Our approach is designed to capitalize on the rampant advances in both natural language modeling and unconditional molecule generation. The adapter is lightweight and compatible with different kinds of pretrained LLM encoders and molecule decoders. Looking forward, the biological and chemical communities have shown strong interest in using natural language to condition the generation of small molecules and even proteins with desired structural and functional properties Pei et al. (2024). ChemLML provides a path for combining the most advanced natural language and domain-specific models as they continue to undergo rapid development for this type of text-conditioned generation.
Acknowledgments
This research was supported by National Institutes of Health award R01GM135631. The research was performed using the computing resources and assistance of the University of Wisconsin-Madison Center for High Throughput Computing (Center for High Throughput Computing, 2006). We thank Shengchao Liu for addressing problems with dataset collection, Yin Fang for discussion of Mol-Instruction, and Chaowei Xiao for insightful discussion about the experimental setting.